# Construction and classification of differential symmetry breaking operators for principal series representations of the pair $(SO_{0}(4,1),SO_{0}(3,1))$ for special parameters

### v1

@@ 4.1 Finite-Dimensional Representation Theory of SO(3)
~$\begin{gathered}\widetilde{{\mathbb{C}}}_{\lambda,\nu}:=\text{{Rest}}_{x_{3}=0}\circ\left(I_{\nu-\lambda}\widetilde{C}^{\lambda-1}_{\nu-\lambda}\right)\left(-\Delta_{{\mathbb{R}}^{2}},\frac{\partial}{\partial x_{3}}\right).\end{gathered}$
@@ 5.2 Reduction Theorems
Now, for $f_{0},f_{\pm 1},f_{\pm 2}\ldots,f_{\pm N}\in{\mathbb{C}}[t]$ set $\mathbf{f}:={}^{t}(f_{-N},\ldots,f_{0},\ldots,f_{N})$ and let $L_{j}^{A,\pm}(\mathbf{f})(t)$ and $L_{j}^{B,\pm}(\mathbf{f})(t)$ $(j=0,\dots,N)$ be the following polynomials:
$\begin{array}[]{l}\begin{rcases*}L_{j}^{A,+}(\mathbf{f})(t):=S_{a+m-j}^{\lambda+j-1}f_{j}-2(N-j)\frac{d}{dt}f_{j+1}.\\[6.0pt] L_{j}^{A,-}(\mathbf{f})(t):=S_{a-m-j}^{\lambda+j-1}f_{-j}+2(N-j)\frac{d}{dt}f_{-j-1}.\end{rcases*}(j=0,1,\ldots,N)\\[30.0pt] \begin{rcases*}L_{j}^{B,+}(\mathbf{f})(t):=2(-m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{j}\\ +(N-j)\frac{d}{dt}f_{j+1}+(N+j)\frac{d}{dt}f_{j-1}.\\[6.0pt] L_{j}^{B,-}(\mathbf{f})(t):=2(m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{-j}\\ -(N+j)\frac{d}{dt}f_{-j+1}-(N-j)\frac{d}{dt}f_{-j-1}.\end{rcases*}(j=1,\ldots,N)\end{array}$ (5.2)
@@
Given $\mathbf{g}=(g_{k_{d}})_{d=0}^{2N}={}^{t}(g_{k_{0}},\ldots,g_{k_{2N}})$ with $g_{k_{d}}\in{\mathbb{C}}[t]$, we write
$\widetilde{\mathbf{g}}={}^{t}(\widetilde{g}_{-N},\ldots,\widetilde{g}_{N}):=\begin{pmatrix}&&1\\ &\iddots&\\ 1&&\end{pmatrix}\mathbf{g}={}^{t}(g_{k_{2N}},\ldots,g_{k_{0}}).$ (5.3)
@@
$M_{s}(\psi)=M_{s}^{\text{{{scalar}}}}(\psi)+M_{s}^{\text{{{vect}}}}(\psi)=0,\enspace\text{ for all }s=0,1,\ldots,2N.$ (6.4)
@@ 6.3 Explicit formulæ for $M_{s}(\psi)$
in terms of the polynomials $L_{j}^{A,\pm}(\widetilde{\mathbf{g}}),L_{j}^{B,\pm}(\widetilde{\mathbf{g}})$ ($j=0,\ldots,N$) defined in (5.2). To do this, we first recall the notion of the $T$-saturation.
@@
[…]
@@
$\begin{gathered}\psi={}^{t}(\psi_{s})_{s=0}^{2N}={}^{t}(\psi_{k-m+N})_{k=m-N}^{m+N}={}^{t}\left(\left(T_{a-k}g_{k}\right)(\zeta_{1}+i\zeta_{2})^{k}\right)_{k=m-N}^{m+N}\\[4.0pt] =\begin{pmatrix}\displaystyle{(T_{a-m+N}g_{m-N})(\zeta)(\zeta_{1}+i\zeta_{2})^{m-N}}\\ \vdots\\ \displaystyle{(T_{a-m}g_{m})(\zeta)(\zeta_{1}+i\zeta_{2})^{m}}\\[4.0pt] \vdots\\ \displaystyle{(T_{a-m-N}g_{m+N})(\zeta)(\zeta_{1}+i\zeta_{2})^{m+N}}\end{pmatrix}.\end{gathered}$
@@
$M_{s}^{\text{{{scalar}}}}(\psi)=$ $\enspace\frac{\zeta_{1}}{Q_{2}(\zeta^{\prime})}T_{a-k}\left(S_{a-k}^{\lambda-1}g_{k}\right)(\zeta_{1}+i\zeta_{2})^{k}$
@@
$R_{k}^{1}(\mathbf{g})=R_{k}^{2}(\mathbf{g})=0\enspace(\forall k=m-N,\dots,m+N)\Leftrightarrow L_{j}^{A,\pm}(\widetilde{\mathbf{g}})=L_{j}^{B,\pm}(\widetilde{\mathbf{g}})=0\enspace(\forall j=0,\dots,N).$
@@
$2R_{k}^{2}(\mathbf{g})-R_{k}^{1}(\mathbf{g})=$ $\enspace S_{a-k}^{\lambda-1}g_{k}+2(m-k)(a-k-\vartheta_{t})g_{k}$
@@
$=$ $\enspace S_{a-k}^{\lambda+m-k-1}g_{k}+4k(\lambda+a-1+m-k)g_{k}-2(N+k-m)\frac{d}{dt}g_{k-1}$
$=$ $\enspace S_{a+k}^{\lambda+m-k-1}g_{k}-2(N+k-m)\frac{d}{dt}g_{k-1},$
which is precisely $L_{j}^{A,+}(\widetilde{\mathbf{g}})$ with $j=m-k=0,1,\ldots,N$. On the other hand, for $k=m,m+1,\ldots,m+N$, and again by (9.8), we have that
@@
which is precisely $L_{j}^{A,-}(\widetilde{\mathbf{g}})$ with $j=k-m=0,1,\ldots,N$.
@@
$R_{k}^{1}(\mathbf{g})-R_{k}^{2}(\mathbf{g})=$ $\enspace 2(-m(\lambda+a-1)+(m-k)(\lambda-1+\vartheta_{t}))g_{k}$
@@
which is $L_{j}^{B,+}(\widetilde{\mathbf{g}})$ with $j=m-k=1,2,\ldots,N$. The identity above is also valid for $k=m+1,m+2,\ldots,m+N$, which is equal to $-L_{j}^{B,-}(\widetilde{\mathbf{g}})$ with $j=k-m=1,2,\ldots,N$. Thus, we have proved the following:
@@
Let $M_{s}\enspace(s=0,\ldots,2N)$ be the vector coefficients defined in (6.10) and let $\mathbf{g}=(g_{k})_{k=m-N}^{m+N}$. Then, the following holds:
@@
$p_{1}:=$ $\enspace T_{a-k-2}\left(S_{a-k}^{\lambda-1}g_{k}+(N-k+m)\frac{d}{dt}g_{k+1}\right),$
$p_{2}:=$ $\enspace T_{a-k-2}\left(2(m-k)(a-k-\vartheta_{t})g_{k}-(N-k+m)\frac{d}{dt}g_{k+1}\right),$
$p_{3}:=$ $\enspace T_{a-k-2}\left(iS_{a-k}^{\lambda-1}g_{k}+2i(k-m)(a-k-\vartheta_{t})g_{k}+2i(N-k+m)\frac{d}{dt}g_{k+1}\right)$
$=$ $\enspace-i(p_{2}-p_{1}),$
$p_{4}:=$ $\enspace T_{a-k}\left(2k(\lambda+a-1+m-k)g_{k}-(N+k-m)\frac{d}{dt}g_{k-1}\right).$
@@
$p_{1}-p_{2}=p_{3}=\enspace Q_{2}(\zeta^{\prime})p_{1}+p_{4}=0,$
@@
The proof is a bit long and uses several technical results, including properties of renormalized Gegenbauer polynomials and imaginary Gegenbauer operators which are listed in the appendix (Section 9). Note that even if it is a linear system of ordinary differential equations with polynomial solutions (apparently easy), it is an overdetermined system, since we have $2(2N+1)$ equations and $12N+$ 1 functions.
@@
    Inductively, for $j=0,\dots,N-1$, obtain $f_{\pm j}$ by solving $(B_{j}^{\pm})$ and by using the expressions of $f_{\pm N}$ obtained in Phase 1. After that, check that these $f_{\pm j}$ satisfy equations $(A_{j}^{\pm})$. For $f_{0}$ we obtain two different expressions (one coming from the $+$ side and another coming from the -- side).
@@ 7.1 Phases 1 and 2: Obtaining $f_{\pm j}$ up to constant
$(A_{N}^{+})\enspace S_{a}^{\lambda+N-1}f_{N}=0,$ $(A_{N}^{-})\enspace S_{a-2N}^{\lambda+N-1}f_{-N}=0.$
@@
Concretely, we write the functions as vertices in a graph (including 0, that can be thought as the starting point), and represent equations $(A_{j}^{\pm})$ and $(B_{j}^{\pm})$ as edges linking the functions that they relate. Equations $(A_{j}^{\pm})$ are written as dotted red lines, while $(B_{j}^{\pm})$ are written as solid blue lines. Moreover, between these edges, we chose a few and write them as arrows, representing the order in which we solve the equations of the system. The first two arrows are dotted and red (corresponding to solving $(A_{N}^{\pm})$ in Phase 1), and the rest are solid and blue (since we obtain the rest of the functions through solving $(B_{j}^{\pm})$ in Phase 2). We remark that in contrast with the dotted red edges (equations $(A_{j}^{\pm})$), the solid blue arrows (equations $(B_{j}^{\pm})$) link three vertices (except for the first one associated to $(B_{N}^{\pm})$, that links only two), as one can deduce from (7.1).
@@
$-(-i)^{N-j}\Gamma_{j}^{+}\frac{d}{dt}f_{j-1}=2q_{N}^{+}\widetilde{C}_{a+j-N-2}^{\lambda+N}(it).$
@@
Now, by using (9.9) for $d=N-j+1$, this amounts to
@@
If $a+j-N$ is even, $c_{j-1}^{+}=0$ necessarily by the condition $f_{j-1}\in\text{{Pol}}_{a-N+j-1}[t]_{\text{even}}$. In particular, (7.9) holds. On the other hand, if $a+j-N$ is odd, from $\Gamma_{j-1}^{+}\neq 0$ we have
$2\gamma(\lambda+N-1,a+j-1-N)=2\lambda+N+j+a-3\neq 0$
@@
for some constant $c_{j-1}^{-}\in{\mathbb{C}}$. Now, by (9.8) for $d=N-j+1$ and (9.6), $(A_{j-1}^{-})$ is easily shown to be equivalent to
@@
As before, if $a-j-N$ is even, $c_{j-1}^{-}=0$ necessarily by the condition $f_{-j+1}\in\text{{Pol}}_{a-N-j+1}[t]_{\text{even}}$. In particular (7.10) is satisfied. On the other hand, if $a-j-N$ is odd, by $\gamma(\lambda,a+j-N-3)\neq 0$ we have
$2\gamma(\lambda,a+j-N-3)=2\lambda+a+j-N-3\neq 0$
@@
$-\frac{d}{dt}f_{N-1}=(\vartheta_{t}-a)f_{N}=2q_{N}^{+}\widetilde{C}_{a-2}^{\lambda+N}(it),$
@@
$-i\Gamma_{N-1}^{+}=-i\gamma(\lambda+N-1,a-1),$
@@
$-i\gamma(\lambda+N-1,a-1)f_{N-1}(t)=q_{N}^{+}\widetilde{C}_{a-1}^{\lambda+N-1}(it)+c_{N-1}^{+},$
@@
which multiplied by $-i\gamma(\lambda+N-1,a-1)\neq 0$ amounts to
@@
$\gamma(\lambda+N-1,a-s)=0,\enspace\text{ for some }s=1,2,\ldots,N.$ (7.14)
@@
Then, in particular, we have $\gamma(\lambda+N-1,a-r)\neq 0$ for any $r=1,2,\ldots,N$, $r\neq s$. By (7.5), this implies that $\Gamma_{j-1}^{+}\neq 0$ for any $j=N-s+2,N-s+3,\ldots,N$. Therefore, by (7.3) and by using Lemma 7.1(1) recursively, we obtain that $f_{j}$ is given by (7.7) for any $j=N-s+1,N-s+2,\ldots,N$ for some constant $q_{N}^{+}\in{\mathbb{C}}$. Now, for $j=N-s$, we have that $(B_{N-s+1}^{+})$ multiplied by $(-i)^{s-1}\Gamma_{N-s+1}^{+}\neq 0$ amounts to
$-(-i)^{s-1}\Gamma_{N-s+1}^{+}\frac{d}{dt}f_{N-s}=2q_{N}^{+}\widetilde{C}_{a-s-1}^{\lambda+N}(it).$
@@
    $(a)_{0}=1\enspace\text{ and }(a)_{\ell}=a(a-1)\cdots(a-\ell+1).$
@@
$\gamma(\lambda+N-1,a-2N-s)=0,\enspace\text{ for some }s=1,2,\ldots,N.$ (7.15)
@@
Hence, it suffices to prove the statement for $j=N-s-n$ when $n=0,1,\ldots,N-s$. We prove this by induction on $n$. Concretely, we prove that a solution of the system must satisfy (7.16), and later, we prove that if $f_{-j}$ for $j=N-s-n$ ($n=0,1,\ldots,N-s$) is given by (7.16), then, $f_{-j}$ provide a solution of the system if and only if (7.17) is satisfied.
@@
Now, suppose that (7.16) holds for $n$ and $n+1$, (i.e., for $f_{s+n-N}$ and $f_{s+n-N-1}$), for some $n=0,\ldots,N-s-1$. Let us show that it also holds for $n-1$ (i.e., for $f_{s+n-N+1}$). For $j=N-s-n,\ldots,N$ we write $f_{-j}$ as
@@
Now, suppose that $f_{-j}$ is given by (7.16) for all $j=0,\ldots,N$. Let us show that $f_{-j}$ provide a solution of the system if and only if the relations (7.17) hold. For $j=N-s+1,\ldots,N$ there is nothing to prove, since $P_{-j}\equiv 0$. Suppose then $j=0,\ldots,N-s$. Since we obtained $f_{-j}$ by solving $(B_{j+1}^{-})$ recursively, it suffices to show that $f_{-j}$ satisfies $(A_{j}^{-})$. Fix $j=N-s-n$ with $n=0,1,\ldots,N$. As pointed out above, $(A_{N-s-n}^{-})$ is satisfied if and only if (7.20) holds. By using (9.10), and the definition of $P_{N-s-n}$, the left-hand side amounts to
@@
Both sides are polynomials of degree $n-2$, so they clearly coincide if and only if each term $t^{n-2\ell-2}$ (for $\ell=0,1,\ldots,\left[\frac{n-2}{2}\right]$) coincide. By a straightforward computation, this is easily shown to by equivalent to (7.17). Hence, the lemma is proved. ∎
Now we consider the case $N<a<2N$. This case can be done by using the same machinery we developed in the case $a\geq 2N$. First, note that Lemma 7.8(1) and Remark 7.4(1) still holds in this case, so there are no changes in the $+$ side. Thus, it suffices to think about the -- side. In this case, instead of distinguishing between $\Gamma_{0}^{-}\neq 0$ and $\Gamma_{0}^{-}=0$, we consider the following two cases:
@@
Since $a-N-j<0$ for $j=a-N+1,a-N+2,\ldots,N$, we have that $f_{-j}\equiv 0$ for $j=a-N+1,a-N+2,\ldots,N$ by the condition $f_{-j}\in\text{{Pol}}_{a-N-j}[t]_{\text{even}}$. Hence, by (7.3) and by using Lemma 7.1(2) recursively for $j=1,2,\ldots,a-N+1$ we can obtain $f_{N-a},\ldots,f_{0}$ as soon as $\gamma(\lambda,a+j-N-3)\neq 0$ for all $j=1,2,\ldots,a-N+1$, which is precisely the condition (a). Hence, in this case we have that $f_{-j}$ is given by (7.13) for all $j=0,1,\ldots,N$.

### v2

@@ 4.1 Finite-Dimensional Representation Theory of $SO(3)$
$\begin{gathered}\widetilde{{\mathbb{C}}}_{\lambda,\nu}:=\text{{Rest}}_{x_{3}=0}\circ\left(I_{\nu-\lambda}\widetilde{C}^{\lambda-1}_{\nu-\lambda}\right)\left(-\Delta_{{\mathbb{R}}^{2}},\frac{\partial}{\partial x_{3}}\right).\end{gathered}$
@@ 5.2 Reduction Theorems
[…]
$\begin{array}[]{l}\begin{rcases*}L_{j}^{A,+}(\mathbf{f})(t):=S_{a+m-j}^{\lambda+j-1}f_{j}-2(N-j)\frac{d}{dt}f_{j+1}.\\[6.0pt] L_{j}^{A,-}(\mathbf{f})(t):=S_{a-m-j}^{\lambda+j-1}f_{-j}+2(N-j)\frac{d}{dt}f_{-j-1}.\end{rcases*}(j=0,1,\ldots,N)\\[30.0pt] \begin{rcases*}L_{j}^{B,+}(\mathbf{f})(t):=2(-m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{j}\\ \hskip 0.0pt plus 1.0+(N-j)\frac{d}{dt}f_{j+1}+(N+j)\frac{d}{dt}f_{j-1}.\\[6.0pt] L_{j}^{B,-}(\mathbf{f})(t):=2(m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{-j}\\ \hskip 0.0pt plus 1.0-(N+j)\frac{d}{dt}f_{-j+1}-(N-j)\frac{d}{dt}f_{-j-1}.\end{rcases*}(j=1,\ldots,N)\end{array}$ (5.2)
@@
[…]
@@
$M_{s}(\psi)=M_{s}^{\text{{{scalar}}}}(\psi)+M_{s}^{\text{{{vect}}}}(\psi)=0,\kern 4.62497pt\text{ for all }s=0,1,\ldots,2N.$ (6.4)
@@ 6.3 Explicit formulæ for $M_{s}(\psi)$
[…]
@@
${\mathbb{C}}[t]$${\mathbb{C}}(\zeta_{1},\zeta_{2},\zeta_{3})$${\mathbb{C}}[t]$${\mathbb{C}}(\zeta_{1},\zeta_{2},\zeta_{3})$$\scriptstyle{T_{\ell}}$$\scriptstyle{S}$$\scriptstyle{D}$$\scriptstyle{T_{\ell}}$
@@
[…]
@@
$M_{s}^{\text{{{scalar}}}}(\psi)=$ $\kern 4.62497pt\frac{\zeta_{1}}{Q_{2}(\zeta^{\prime})}T_{a-k}\left(S_{a-k}^{\lambda-1}g_{k}\right)(\zeta_{1}+i\zeta_{2})^{k}$
@@
$R_{k}^{1}(\mathbf{g})=R_{k}^{2}(\mathbf{g})=0\kern 4.62497pt(\forall k=m-N,\dots,m+N)\Leftrightarrow L_{j}^{A,\pm}(\widetilde{\mathbf{g}})=L_{j}^{B,\pm}(\widetilde{\mathbf{g}})=0\kern 4.62497pt(\forall j=0,\dots,N).$
@@
$2R_{k}^{2}(\mathbf{g})-R_{k}^{1}(\mathbf{g})=$ $\kern 4.62497ptS_{a-k}^{\lambda-1}g_{k}+2(m-k)(a-k-\vartheta_{t})g_{k}$
@@
$=$ $\kern 4.62497ptS_{a-k}^{\lambda+m-k-1}g_{k}+4k(\lambda+a-1+m-k)g_{k}-2(N+k-m)\frac{d}{dt}g_{k-1}$
$=$ $\kern 4.62497ptS_{a+k}^{\lambda+m-k-1}g_{k}-2(N+k-m)\frac{d}{dt}g_{k-1},$
[…]
@@
[…]
@@
$R_{k}^{1}(\mathbf{g})-R_{k}^{2}(\mathbf{g})=$ $\kern 4.62497pt2(-m(\lambda+a-1)+(m-k)(\lambda-1+\vartheta_{t}))g_{k}$
@@
[…]
@@
Let $M_{s}\kern 4.62497pt(s=0,\ldots,2N)$ be the vector coefficients defined in (6.10) and let $\mathbf{g}=(g_{k})_{k=m-N}^{m+N}$. Then, the following holds:
@@
$p_{1}$= $\kern 4.62497ptT_{a-k-2}\left(S_{a-k}^{\lambda-1}g_{k}+(N-k+m)\frac{d}{dt}g_{k+1}\right),$
$p_{2}$= $\kern 4.62497ptT_{a-k-2}\left(2(m-k)(a-k-\vartheta_{t})g_{k}-(N-k+m)\frac{d}{dt}g_{k+1}\right),$
$p_{3}$= $\kern 4.62497ptT_{a-k-2}\left(iS_{a-k}^{\lambda-1}g_{k}+2i(k-m)(a-k-\vartheta_{t})g_{k}+2i(N-k+m)\frac{d}{dt}g_{k+1}\right)$
$=$ $\kern 4.62497pt-i(p_{2}-p_{1}),$
$p_{4}$= $\kern 4.62497ptT_{a-k}\left(2k(\lambda+a-1+m-k)g_{k}-(N+k-m)\frac{d}{dt}g_{k-1}\right).$
@@
$p_{1}-p_{2}=p_{3}=\kern 4.62497ptQ_{2}(\zeta^{\prime})p_{1}+p_{4}=0,$
@@
The proof is a bit long and uses several technical results, including properties of renormalized Gegenbauer polynomials and imaginary Gegenbauer operators which are listed in the appendix (Section 9). Note that even if it is a linear system of ordinary differential equations with polynomial solutions (apparently easy), it is an overdetermined system, since we have $2(2N+1)$ equations and $2N+1$ functions.
@@
    Inductively, for $j=0,\dots,N-1$, obtain $f_{\pm j}$ by solving $(B_{j}^{\pm})$ and by using the expressions of $f_{\pm N}$ obtained in Phase 1. After that, check that these $f_{\pm j}$ satisfy equations $(A_{j}^{\pm})$. For $f_{0}$ we obtain two different expressions (one coming from the ++ side and another coming from the - side).
@@ 7.1 Phases 1 and 2: Obtaining $f_{\pm j}$ up to constant
$(A_{N}^{+})\kern 4.62497ptS_{a}^{\lambda+N-1}f_{N}=0,$ $(A_{N}^{-})\kern 4.62497ptS_{a-2N}^{\lambda+N-1}f_{-N}=0.$
@@
Concretely, we write the functions as vertices in a graph (including 0, that can be thought of as the starting point), and represent equations $(A_{j}^{\pm})$ and $(B_{j}^{\pm})$ as edges linking the functions that they relate. Equations $(A_{j}^{\pm})$ are written as dotted red lines, while $(B_{j}^{\pm})$ are written as solid blue lines. Moreover, between these edges, we chose a few and write them as arrows, representing the order in which we solve the equations of the system. The first two arrows are dotted and red (corresponding to solving $(A_{N}^{\pm})$ in Phase 1), and the rest are solid and blue (since we obtain the rest of the functions through solving $(B_{j}^{\pm})$ in Phase 2). We remark that in contrast with the dotted red edges (equations $(A_{j}^{\pm})$), the solid blue arrows (equations $(B_{j}^{\pm})$) link three vertices (except for the first one associated to $(B_{N}^{\pm})$, that links only two), as one can deduce from (7.1).
@@
$(-i)^{N-j}\Gamma_{j}^{+}\frac{d}{dt}f_{j-1}=2q_{N}^{+}\widetilde{C}_{a+j-N-2}^{\lambda+N}(it).$
@@
Now, by using (9.9) for d=Nj+1d=N-j+1, this amounts to
@@
If a+jNa+j-N is even, $c_{j-1}^{+}=0$ necessarily by the condition $f_{j-1}\in\text{{Pol}}_{a-N+j-1}[t]_{\text{even}}$. In particular, (7.9) holds. On the other hand, if a+jNa+j-N is odd, from $\Gamma_{j-1}^{+}\neq 0$ we have
$2\gamma(\lambda+N-1,a+j-1-N)=2\lambda+N+j+a-3\neq 0.$
@@
for some constant $c_{j-1}^{-}\in{\mathbb{C}}$. Now, by (9.8) for d=Nj+1d=N-j+1 and (9.6), $(A_{j-1}^{-})$ is easily shown to be equivalent to
@@
As before, if ajNa-j-N is even, $c_{j-1}^{-}=0$ necessarily by the condition $f_{-j+1}\in\text{{Pol}}_{a-N-j+1}[t]_{\text{even}}$. In particular (7.10) is satisfied. On the other hand, if ajNa-j-N is odd, by $\gamma(\lambda,a+j-N-3)\neq 0$ we have
$2\gamma(\lambda,a+j-N-3)=2\lambda+a+j-N-3\neq 0.$
@@
$\frac{d}{dt}f_{N-1}=(\vartheta_{t}-a)f_{N}=2q_{N}^{+}\widetilde{C}_{a-2}^{\lambda+N}(it),$
@@
$i\Gamma_{N-1}^{+}=-i\gamma(\lambda+N-1,a-1),$
@@
$i\gamma(\lambda+N-1,a-1)f_{N-1}(t)=q_{N}^{+}\widetilde{C}_{a-1}^{\lambda+N-1}(it)+c_{N-1}^{+},$
@@
which multiplied by $i\gamma(\lambda+N-1,a-1)\neq 0$ amounts to
@@
$\gamma(\lambda+N-1,a-s)=0,\kern 4.62497pt\text{ for some }s=1,2,\ldots,N.$ (7.14)
@@
Then, in particular, we have $\gamma(\lambda+N-1,a-r)\neq 0$ for any $r=1,2,\ldots,N$, $r\neq s$. By (7.5), this implies that $\Gamma_{j-1}^{+}\neq 0$ for any $j=N-s+2,N-s+3,\ldots,N$. Therefore, by (7.3) and by using Lemma 7.1(1) recursively, we obtain that $f_{j}$ is given by (7.7) for any $j=N-s+1,N-s+2,\ldots,N$ for some constant $q_{N}^{+}\in{\mathbb{C}}$. Now, for j=Nsj=N-s, we have that $(B_{N-s+1}^{+})$ multiplied by $(-i)^{s-1}\Gamma_{N-s+1}^{+}\neq 0$ amounts to
$(-i)^{s-1}\Gamma_{N-s+1}^{+}\frac{d}{dt}f_{N-s}=2q_{N}^{+}\widetilde{C}_{a-s-1}^{\lambda+N}(it).$
@@
    $(a)_{0}=1\kern 4.62497pt\text{ and }(a)_{\ell}=a(a-1)\cdots(a-\ell+1).$
@@
$\gamma(\lambda+N-1,a-2N-s)=0,\kern 4.62497pt\text{ for some }s=1,2,\ldots,N.$ (7.15)
@@
Hence, it suffices to prove the statement for j=Nsnj=N-s-n when $n=0,1,\ldots,N-s$. We prove this by induction on $n$. Concretely, we prove that a solution of the system must satisfy (7.16), and later, we prove that if $f_{-j}$ for j=Nsnj=N-s-n ($n=0,1,\ldots,N-s$) is given by (7.16), then, $f_{-j}$ provide a solution of the system if and only if (7.17) is satisfied.
@@
Now, suppose that (7.16) holds for $n$ and $n+1$, (i.e., for $f_{s+n-N}$ and $f_{s+n-N-1}$), for some $n=0,\ldots,N-s-1$. Let us show that it also holds for n1n-1 (i.e., for $f_{s+n-N+1}$). For $j=N-s-n,\ldots,N$ we write $f_{-j}$ as
@@
Now, suppose that $f_{-j}$ is given by (7.16) for all $j=0,\ldots,N$. Let us show that $f_{-j}$ provide a solution of the system if and only if the relations (7.17) hold. For $j=N-s+1,\ldots,N$ there is nothing to prove, since $P_{-j}\equiv 0$. Suppose then $j=0,\ldots,N-s$. Since we obtained $f_{-j}$ by solving $(B_{j+1}^{-})$ recursively, it suffices to show that $f_{-j}$ satisfies $(A_{j}^{-})$. Fix j=Nsnj=N-s-n with $n=0,1,\ldots,N$. As pointed out above, $(A_{N-s-n}^{-})$ is satisfied if and only if (7.20) holds. By using (9.10), and the definition of $P_{N-s-n}$, the left-hand side amounts to
@@
Both sides are polynomials of degree n2n-2, so they clearly coincide if and only if each term $t^{n-2\ell-2}$ (for $\ell=0,1,\ldots,\left[\frac{n-2}{2}\right]$) coincide. By a straightforward computation, this is easily shown to by equivalent to (7.17). Hence, the lemma is proved. ∎
Now we consider the case $N<a<2N$. This case can be done by using the same machinery we developed in the case $a\geq 2N$. First, note that Lemma 7.8(1) and Remark 7.4(1) still holds in this case, so there are no changes in the ++ side. Thus, it suffices to think about the - side. In this case, instead of distinguishing between $\Gamma_{0}^{-}\neq 0$ and $\Gamma_{0}^{-}=0$, we consider the following two cases:
@@
Since aNj<0a-N-j<0 for $j=a-N+1,a-N+2,\ldots,N$, we have that $f_{-j}\equiv 0$ for $j=a-N+1,a-N+2,\ldots,N$ by the condition $f_{-j}\in\text{{Pol}}_{a-N-j}[t]_{\text{even}}$. Hence, by (7.3) and by using Lemma 7.1(2) recursively for $j=1,2,\ldots,a-N+1$ we can obtain $f_{N-a},\ldots,f_{0}$ as soon as $\gamma(\lambda,a+j-N-3)\neq 0$ for all $j=1,2,\ldots,a-N+1$, which is precisely the condition (a). Hence, in this case we have that $f_{-j}$ is given by (7.13) for all $j=0,1,\ldots,N$.